OpenStudy (anonymous):

–2(x – 3) ≥ 5 – (x + 3) can someone show me how to solve this step by step ? i know the answer is x less than or equal to 4 .

OpenStudy (anonymous):

@whpalmer4 @whydoihavetosignupl @EffImHighOffSkittles @eaileen @R00nnYBraiNsbiG @terenzreignz @uri @IloveYou* @primeralph @prettyboy20202 @ajprincess @AnElephant @SnuggieLad @snorl4x @Directrix @dan815 @divagirl421 @DarkKat @Dmitt @faix95 @GOODMAN @hobbs978 @jhonyy9 @Jack17 @KawasumiKimiko @Luigi0210 @Loujoelou @love_jessika15 @lkkk @Callisto @countonme123 @ChrisJP @nincompoop @mukushla @marylou004 @mlb004 @MissHN182 @thefon19 @ryanjnfish @ladycomical @DDCamp @ryanjnfish @jd1028 @phantomofthemathwork @jtrich13 @ajprincess @PaulaCamacho @laxman03 @beccabee77 @babyslapmafro @whpalmer4 @jtrich13 @resident2013 @sumi29 @R00nnYBraiNsbiG @tfuller @Frostbite @SnuggieLad @raque766 @physicsboffin @Falco276 @e.cociuba @blurbendy @kausarsalley @sakigirl @RANE @Danny_Boy @SomeGirl1999 @Jamierox4ev3r @iforgot @lilkittykat36 @n14r96 @ChanceBBrick @knowel @meldo16 @Helpmeplease99 @Manda_luvsu @baseballer20142 @jhalt @RIUKUS @harpelletjain @Math81 @Frank @WildKat @maurimath @shamim @AliaKnight @MayMay_69 @amyamanda @victoria650 @Social @Whatever1 @xxAshxx @lurrch @djoutten @musicalcat98 @Obviously717 @ihategeometry2425 @Zoe_Bailey @RIUKUS @henendez @shaneetta @niah2411 @harpelletjain @Math81 @WildKat @eidan @Hassan07 @Asuman @Frank @bani2100 @turtlebuddy @maurimath @lizzy89 @dborder @peush @EricaaaaB @yehonatan @tori45 @Ojasmin @Markbruce @april115 @RH @shamim @Ryaan @emilyhaddad @Jamierox4ev3r @RANE @NoelGreco @shkrina @lakenclark @Fazeelayaz @sarahusher @JazzyFrizo16 @music101 @HaperFink22 @whydoihavetosignup1 @domu @ashleyjade @Gikundab @Will.Neko @QuinterrianW_23 @tillman--- @baseballer20142 @Manda_luvsu @Charollete @amyamanda @harpreetsk @Whatever1 @HaydenBurden1996 @xxAshxx @lurrch @djoutten @musicalcat98 @Obviously717 @ihategeometry2425 @shaneetta @DLS @NoelGreco @emilyhaddad @marylou004 @JazzyFrizo16 @music101 @HaperFink22 @Courage. @KingAditya @bradely @Gikundab @Will.Neko @QuinterrianW_23 @mcordray @_anastasia0 @ozhobbits @gabigurl @musicalcat98 @Obviously717 @annemoreno @glamourariss @jbsharpe @WildKat @turtlebuddy @turtlebuddy @peush @yehonatan @tori45 @debrajane @EricaaaaB @linabeanax3 @Ojasmin @Markbruce @confusao @sircrashalot @iizabell @tori45 @yehonatan @peush @maurimath @jbsharpe @glamourariss @annemoreno @Obviously717 @mcordray @_anastasia0 @AppylPye @gabigurl @Imtiaz7 @fringe @khadeeja @jhalt @itsmeeQUENNIE @KingAditya @sambhav__jain @whydoihavetosignup1 @mlb004 @Yasminelove1 @caozeyuan @marylou004 @genius12 @agent0smith @AnElephant @lakenclark @JazzyFrizo16 @HaperFink22 @music101 @danielle02 @HaydenBurden1996 @xxAshxx @Obviously717 @Zoe-bailey @peush @yehonatan @tori45 @debrajane @EricaaaaB @linabeanax3 @Fabien1 @Ojasmin @sircrashalot @ankit042 @caozeyuan @rperez36 @freddy_eighty7 @Charollete @freddy_eighty7 @toolCoolChris @agent0smith @genius12 @uri @walac @Luigi0210 @kelliegirl33 @nincompoop @mathmate @stgreen @henpen @binary3i @ChrisJP @zpupster @ankit042 @caozeyuan @betterwiththelightsoff99 @Yasminelove1 @HaperFink22 @glitterythings @whydoihavetosignup1 @Charollete @bradely @ashleyjade @Alexis2332 @Will.Neko @walac @Gikundab @Imtiaz7 @Dianatorres @Alyssa_May_De_Jesus @baseballer20142 @victoria650 @lakenclark @zpupster @mathmate @binary3i @austinL @Loser66 @GoldPhenoix @e.cociuba @DLS @razor99 @Mashy @Machida @RANE @Goten77 @lncognlto @Fifciol @ivancsc1996 @kenbaby @kryton1212 @ankit042 @Henry.Lister @JazzyFrizo16 @betterwiththelightsoff99 @mlb004 @ivano_logy @xxkerstie.bbyxx @MayMay_69 @KawasumiKimiko @itsmeeQUENNIE @Shadly @Albany_Goon @10thgrader @xxmarcusxx1994 @jessy1289 @baseballer20142 @groges @pdd21 @claringleek @Rowa @daniellecrook @imcool @harpreetsk @angelj @JJB @OpenSessame @xseahaven @Whatever1 @harliedanielfrye @borimexz @Jdubose2 @BrianaShamwell @freddy_eighty7 @Horush @snewton29 @Fabien1 @Peter_szlapak @rmpjingwei @mtalampas21 @andriani @mygumshurt @lamel1993 @LittleAnnn @psychomath @yehonatan @heights @JupitersBrain @slayer00647 @2old4math @munastar @jmensal @TanteAnna @Matt-Teepell @DJova @quietkat @bcrisco @larlar77 @Markevia @utvet @felixt03 @9995950976 @natyxz @lftyoutofthepen @kiah44 @Henry.Lister @helder_edwin @Loser66 @inkyvoyd @mathmate @hartnn @dmezzullo @abb0t @amistre64

OpenStudy (falco276):

DOES SOMEONE REALIZE THAT THESE ARE HACKERS?????? CONSTANTLY TAGGING PEOPLE OVER AND OVER?????? TYPE @ AND THE USERS NAME, IT WON'T SHOW UP.

OpenStudy (anonymous):

first we will distribute through the parenthesis -2(x - 3) >= 5 - (x + 3) -2x + 6 >= 5 -x - 3 combine like terms -2x + 6 >= 2 - x add x to both side -2x + x + 6 >= 2 - x + x combine like terms -x + 6 >= 2 now subtract 6 from each side -x + 6 - 6 >= 2 - 6 combine like terms -x >= -4 (now we will multiply by -1 to make x positive. Keep in mind, when multiplying/dividing by a negative, changes the inequality sign. x <= 4 there it is...step by step.....do you have any questions ?